Your math teacher says you have an 86.5% test average after 8 tests. You will be taking 4 more tests and hope to raise your overall test average to 90%. What do you need to average on the next 4 tests in order to reach that goal?
Found 2 solutions by Fombitz, checkley77:
Answer by Fombitz(32388)

The average of your first 8 scores, A, is the sum of the 8 scores divided by 8.
The average of your 12 scores,B, is the sum of the 12 scores divided by 12.
A=SUM8/8
SUM8=8A
SUM8=8(0.865)
SUM8=6.92
B=SUM12/12
SUM12=12B
SUM12=12(.90)
SUM12=10.8
SUM12=SUM8+NINE+TEN+ELEVEN+TWELVE
where NINE, TEN, ELEVEN, and TWELVE are test scores 9,10,11, and 12.
SUM12=SUM8+NINE+TEN+ELEVEN+TWELVE
10.8=6.92+NINE+TEN+ELEVEN+TWELVE
NINE+TEN+ELEVEN+TWELVE=3.88
(NINE+TEN+ELEVEN+TWELVE)/4=3.88/4
(NINE+TEN+ELEVEN+TWELVE)/4=0.97
You need to average a 97% on the next four tests to get 90% average.

Answer by checkley77(12844)

(8*86.5+4x)/12=90
(692+4x)/12=90
692+4x=12*90
4x=-692+1080
4x=388
x=388/2
x=97 is the required average for the remaining 4 test scores.